OpenStudy (anonymous):

Simplify (12z^2-25z+12)/(3z^2+2z-8)

OpenStudy (campbell_st):

factorise the denominator and numerator 12z^2 - 25z +12 = (3z -4)(4z -3) and then factorise 3z^2 +2z -8 = (z +2)(3z -4) cancel common factors... (3z - 4) leaves (4z -3)/(z +2)